OpenStudy (anonymous):

Find the missing coefficient. (4d-4)(3d-2)=12d^2+blank d+8

OpenStudy (midhun.madhu1987):

LHS = (4d-4)(3d-2) = (4d)(3d) + (4d)(-2) + (-4)(3d) + (-4)(-2) = 12d^2 – 8d – 12d +8 = 12d^2 – 20d + 8 ------ (1) But RHS = 12d^2 + _d + 8 ---------(2) Therefore from (1) and (2), Missing Coefficient is -20.